News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 1 December 2000 10.30am EST

Last week the observing schedule was interrupted following a high radiation event at 2:46pm EST on Nov 26 that stopped the daily load and activated the science instrument safing sequence (SCS 107). All actions were nominal and assessment of the dump data confirmed that the trigger was due to high proton counts in the EPHIN P4GM channel exceeding the pre-defined threshold of 300 counts/cm^2/s/sr. The solar levels subsided to an acceptable level over the next 36 hours and the loads were replanned and resumed at 5am EST on 28 Nov. The observations of K284, M31, N103B were impacted and will be rescheduled in a future load.

Real-time procedures were completed during last week to re-enable SCS 107 and radmon on 26 Nov in preparation for load re-start, perform an ACIS CTI measurement on 26 Nov, perform a dump of the ACA software patch table on 28 Nov to verify operation of the ACA fid light software patch, and to reinitialize the OFP thruster command processing subfunction following a counter roll-over. All procedures were completed nominally.

The planned observations for the coming week are as follows and real-time procedures will be conducted to insert and retract the LETG on Dec 5 and 7 respectively.

------------------------------------------
NGC 1068                ACIS-S/HETG  Dec 4  
NGC 1068                ACIS-S/LETG  Dec 5  
PKS2155-304             ACIS-S/LETG  Dec 6  
Radiation Belt
PKS2155-304             ACIS-S/LETG  Dec 7  
AGX J035048-1147        ACIS-I
SNR 1987A               ACIS-S
3C129                   ACIS-S       Dec 8  
Radiation Belt                       Dec 9  
1150+497                ACIS-S
E0102-72                ACIS-S       Dec 10

------------------------------------------

All spacecraft subsystems continued to operate nominally.

Return to Status Reports